OpenStudy (anonymous):

given log 3 = 0.48 and log 7 = 0.84 calculate the value of log (0.30/0.7)

OpenStudy (anonymous):

here's a hint that might help... \[\huge \log a - \log b = \log (\frac ab )\] does that give you any ideas?

OpenStudy (anonymous):

uhm...what's huge and what's frac ab?

OpenStudy (anonymous):

refresh your screen. you can't seem to see the latex

OpenStudy (anonymous):

what's written there is log a - log b = log (a/b)

OpenStudy (anonymous):

let me give an example log 6 = 2 log 4 = 3 then i want to find log (6/4) so i do: log (6/4) = log 6 - log 4 so if i substitute the values i have... log (6/4) = 2 - 3 so log (6/4) = -1 do you get the idea?

OpenStudy (anonymous):

yeah, i know that, but then its 0.30 n 0.7....so what bout that?

OpenStudy (anonymous):

do you agree that 0.30 is equal to 3/10 and 0.7 is equal to 7/10 ?

OpenStudy (anonymous):

yes....i tried that, but still cudnt get it :(

OpenStudy (anonymous):

so you know that 0.3 is 3/10 and 0.7 is 7/10 now here comes the magic 0.3/0.7 is equal to (3/10) / (7/10) if you remember your algebra, you know that the denominator 10 would just cancel out. so you have 0.3/0.7 is equal to 3/7

OpenStudy (anonymous):

but then u cant just cancel out 10..in logarithms i mean..

OpenStudy (anonymous):

its just log3 - log7

OpenStudy (anonymous):

u could

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

hmm. we;; there would be another way to approach this log (0.3) = log (3/10) = log 3 - log 10 log (0.7) = log (7/10) = log 7 - log 10 get the value of those two then find the difference of those differences

OpenStudy (anonymous):

so,... (0.48 - 1) - (0.84 - 1) so its bar1 . 48 - bar1.84 right?

OpenStudy (anonymous):

yes..although i do not know what you mean by bar1

OpenStudy (anonymous):

minus one....

OpenStudy (anonymous):

the minus on top of the 1 - 1

OpenStudy (anonymous):

so is the answer bar1.64?

OpenStudy (anonymous):

let me try

OpenStudy (anonymous):

unfortunately no. what did you do?

OpenStudy (anonymous):

huh? what's the answer?

OpenStudy (anonymous):

the answer should be negative.. since you're doing 0.48 - 1 and 0.84 - 1

OpenStudy (anonymous):

ya then?

OpenStudy (anonymous):

thats wat i said... negative one point 64

OpenStudy (anonymous):

oh. bar means negative. anyway.. -1.64 is still not the answer. 0.48 - 1 is -0.52 0.84 - 1 is -0.36 subtract those two. remember that subtracting negative becomes positive (i.e. a - (-b) = a + b)

OpenStudy (anonymous):

but there isnt a minus in logs...the minus either applies only for the characteristic or mantissa

OpenStudy (anonymous):

who said logs cant have negatives? let me give you an example 2^(-1) = 0.5 if you put that in log form log_2 (0.5) = -1

OpenStudy (anonymous):

the only time you can't have negatives is when you have log (-x) that's illegal since the base of that log is even and there's no exponent that can make an even number into a negative number

OpenStudy (mathstudent55):

log(0.3/0.7) = log(3/7) = log3 - log7 = 0.48 - 0.84 = -0.36 log(0.3/0.7) = log[ 3/10) / (7/10) ] = log(3/10) - log(7/10) = log3 - log10 - (log 7 - log10) = log3 - log10 - log7 + log10 = log 3 - log7 + log10 - log10 = log3 - log7 = 0.48 - 0.84 = -0.36

OpenStudy (anonymous):

well im getting the answer as bar1 point 64

OpenStudy (mathstudent55):

bar 1 point 64 means -1 + 0.64 = -0.36, so you are correct

OpenStudy (anonymous):

Oh.. :) thanks

OpenStudy (mathstudent55):

wlcm